### v2

# ImAgent: A Unified Multimodal Agent Framework for Test-Time Scalable Image Generation

Abstract
Recent text-to-image (T2I) models have made remarkable progress in generating visually realistic and semantically coherent images. However, they still suffer from randomness and inconsistency with the given prompts, particularly when textual descriptions are vague or underspecified. Existing approaches, such as prompt rewriting, best-of-N sampling, and self-refinement, can mitigate these issues but usually require additional modules and operate independently, hindering test-time scaling efficiency and increasing computational overhead. In this paper, we introduce ImAgent, a training-free unified multimodal agent that integrates reasoning, generation, and self-evaluation within a single framework for efficient test-time scaling. Guided by a policy controller, multiple generation actions dynamically interact and self-organize to enhance image fidelity and semantic alignment without relying on external models. Extensive experiments on image generation and editing tasks demonstrate that ImAgent consistently improves over the backbone and even surpasses other strong baselines where the backbone model fails, highlighting the potential of unified multimodal agents for adaptive and efficient image generation under test-time scaling.
1 Introduction
Text-to-Image (T2I) models [rombach2022high, ramesh2022hierarchical, podell2023sdxl, sun2024autoregressive, tian2024visual, qin2025lumina] have made remarkable progress in generating visually realistic and semantically coherent images from natural language descriptions. Despite these advancements, the quality of generated results often exhibits randomness and inconsistency with the given prompts [xu2024prompt, lee2024direct]. This limitation primarily stems from the strong dependency of existing T2I models on the clarity and specificity of textual inputs. When the prompt is vague or underspecified, the models tend to overlook key semantic elements, leading to images that deviate from the intended meaning and fail to fully capture the user’s intent.
To mitigate this problem, prior research has proposed various strategies, such as prompt rewriting [brade2023promptify, mo2024dynamic, liu2025one], best-of-N sampling [li2025reflect, singhal2025general], classifier-free guidance [ho2022classifier, sauer2023stylegan, shen2024rethinking], and self-revision or iterative refinement [singh2023divide, li2024g, jeon2025iterative]. These approaches aim to either clarify the textual prompt or reduce the model’s inherent randomness rather than finetuning the models, reflecting the underlying principle of test-time scaling—improving generation quality through additional inference-time computation.
While effective, these methods typically require additional components, such as a language model for prompt enhancement, a generative model for image synthesis, and a vision-language model for evaluation, thereby increasing memory consumption and computational cost. Moreover, they are often applied independently, requiring human intervention to determine the optimal approach for a given case, which substantially limits the efficiency of test-time scaling and leads to unnecessary computational overhead.
In this paper, we aim to construct a universal agent for image generation that can adaptively select the optimal action for a given case, allocate computational resources accordingly, and execute the selected action within the agent itself without relying on any external models. This design enables more efficient test-time scaling. Thanks to recent advances in unified multimodal models [deng2025emerging, wu2025janus, wu2025janus, xie2024show, qin2025lumina], which integrate text generation, image generation, and visual understanding within a single framework, this assumption becomes attainable. Building upon this foundation, we introduce ImAgent, a training-free unified multimodal agent designed to perform efficient test-time scaling for image generation.
Specifically, ImAgent is built around a policy controller that serves as its “brain”, determining whether an image requires refinement based on the observation history and deciding which action should be invoked in the next step. Under its coordination, multiple predefined generation actions, such as Prompt Enhancement with CoT, Image Detail Refinement, and Best-of-N Sampling, operate collaboratively within a unified framework. Notably, all these actions are executed internally through the agent itself. Through this dynamic interaction, ImAgent transforms what was previously a manually assembled pipeline into a self-organizing agent capable of reasoning, generation, and self-improvement, thereby enhancing the efficiency of test-time scaling.
We conduct experiments on two popular tasks, image generation and image editing, to evaluate the effectiveness of the proposed ImAgent. Extensive experimental results show that ImAgent achieves significant improvements over the backbone model, demonstrating its strong capability and the effectiveness of test-time scaling in enhancing image generation performance. The contributions of this paper are summarized as follows:
We propose a unified multimodal agent framework for image generation, ImAgent, which integrates multiple generation actions and dynamically coordinates them through a policy controller.
Extensive experiments on image generation and editing tasks demonstrate that ImAgent achieves substantial improvements over the backbone model, even outperforming baselines that the backbone model fails to surpass, validating its effectiveness and in test-time scaling.
2 Related Work
2.1 Text-to-Image Generation
Text-to-Image (T2I) models have demonstrated remarkable ability in synthesizing visually realistic and semantically coherent images from natural language descriptions. The evolution of T2I models has followed several major paradigms, including diffusion-based models [podell2023sdxl, rombach2022high, ramesh2022hierarchical, esser2024scaling, gao2024lumina, xie2025sana], autoregressive models [sun2024autoregressive, zhang2024var, tian2024visual, wang2024emu3, zheng2025parallel, xiong2025llava], and more recently, unified multimodal models [xiao2025omnigen, xie2024show, zhou2024transfusion, tong2025metamorph, deng2025emerging, zheng2025learning, luo2024stable]. These approaches have significantly advanced image fidelity, diversity, and text-image alignment.
2.2 Unified Multimodal Models
Unified multimodal models refer to models that jointly possess multimodal understanding and generation capabilities within a single framework. Unlike traditional pipelines that rely on separate models for text or image understanding, image generation, and image editing, these models integrate vision and language processing into a shared architecture, enabling seamless information exchange across modalities [xiao2025omnigen, zhou2024transfusion, tong2025metamorph, wu2025janus, wu2025janus]. This convergence of understanding and generation not only improves efficiency and coherence but also paves the way for building more general-purpose, human-like agents capable of reasoning, creating, and self-correcting.
2.3 Generation Optimization Strategies
Despite the remarkable progress of T2I models, the quality of generated images still depends heavily on the clarity of prompts and the randomness inherent in the generation process. To alleviate these issues, various optimization strategies have been proposed. Best-of-N sampling [tian2025unigen, verdun2025soft, li2025reflect] generates multiple candidates and selects the best one, reducing the influence of stochastic variations in model sampling. Prompt rewriting [tian2025unigen, wang2024promptcharm, yasunaga2022retrieval] reformulates or enriches the input text to provide more explicit guidance for the generator, improving semantic alignment and visual coherence. classifier-free guidance [ho2022classifier, sauer2023stylegan, shen2024rethinking] controls the conditioning strength to balance fidelity and diversity, and iterative self-refinement [yang2023idea2img, madaan2023self, singh2023divide, li2024g, jeon2025iterative] enables models to assess and revise their outputs through multiple rounds of generation. Although these techniques significantly enhance performance, they typically operate as independent modules with limited interaction between reasoning, generation, and evaluation—highlighting the need for unified multimodal frameworks that integrate these capabilities into a cohesive pipeline.
3 Method
We construct ImAgent upon unified multimodal models, which inherently possess both understanding and generation capabilities. Motivated by the observation that multimodal understanding is generally more reliable and easier to achieve than multimodal generation [zhang2025unified, zhang2025unified1], ImAgent leverages the model’s strong understanding ability to adaptively select the optimal action for a given case that enhances the overall generation quality. This design enables ImAgent to fully exploit the potential of a single unified multimodal model without any additional training or external modules, thereby achieving efficient test-time scaling. In this section, we first introduce the overall agent framework of ImAgent, followed by a detailed description of its action space.
3.1 Agent Framework Overview
As shown in Figure 2, ImAgent dynamically selects the next action based on the current observation and executes the selected action until either the STOP action is chosen or the maximum step limit is reached. Formally, we define the state at the current -th time step as:
where and denote the initial user prompt and input image, and represent the current prompt and generated image at -th step, and denotes the history of past observations, i.e., action history.
Then, ImAgent dynamically selects an action from the defined action space (detailed in Section 3.2) using a policy controller conditioned on the current state at -th step:
Upon selecting , the corresponding action function is invoked to produce an updated prompt , an updated image , and a new observation :
Then, the observation is updated as follows:
Unlike conventional agent-based frameworks where observations directly influence state updates, in ImAgent the observation history is primarily used for action selection, while prompt and image updates are determined by the selected generation actions.
This iterative process continues until the agent selects the STOP action, which indicates that the generated result is satisfactory, or until the maximum number of steps (set to 5 by default) is reached. Notably, both the policy controller and the action execution are implemented within a single unified multimodal model. The detailed algorithm for image generation is presented in Algorithm 1, while the algorithm for image editing follows a similar procedure and is provided in the Appendix.
| Types | Model | Comm. | Comp. | Logical | Numerical | Causal | Overall |
| SD3-medium | 0.54 | 0.64 | 0.55 | 0.50 | 0.18 | 0.53 | |
| Sana-1.5 | 0.49 | 0.67 | 0.49 | 0.48 | 0.21 | 0.49 | |
| Lumina-T2I | 0.38 | 0.49 | 0.38 | 0.45 | 0.18 | 0.39 | |
| Omnigen | 0.43 | 0.60 | 0.51 | 0.47 | 0.34 | 0.48 | |
| LLM4GENSD1.5 | 0.55 | 0.48 | 0.55 | 0.39 | 0.45 | 0.51 | |
| ELLASD1.5 | 0.40 | 0.44 | 0.40 | 0.32 | 0.29 | 0.39 | |
| LlamaGen | 0.38 | 0.39 | 0.38 | 0.35 | 0.12 | 0.36 | |
| DALL-E-3 | 0.78 | 0.76 | 0.69 | 0.69 | 0.64 | 0.73 | |
| Gen | gpt-image-1 | 0.83 | 0.87 | 0.81 | 0.88 | 0.71 | 0.83 |
| EMU3 | 0.44 | 0.59 | 0.55 | 0.61 | 0.41 | 0.52 | |
| Show-o | 0.42 | 0.59 | 0.42 | 0.57 | 0.30 | 0.46 | |
| Unified | Lumina-Image 2.0 | 0.49 | 0.65 | 0.56 | 0.43 | 0.40 | 0.52 |
| Vanilla | 0.47 | 0.65 | 0.53 | 0.68 | 0.40 | 0.54 | |
| Bagel | ImAgent | 0.58 | 0.68 | 0.62 | 0.71 | 0.53 | 0.62 |
| Vanilla | 0.45 | 0.60 | 0.46 | 0.46 | 0.36 | 0.47 | |
| Janus-Pro-7B | ImAgent | 0.47 | 0.58 | 0.52 | 0.54 | 0.44 | 0.51 |
3.2 Action Space
In this section, we introduce the action space of ImAgent. The detailed definitions and usage of each action are provided in the Appendix.
Naive Generation/Editing.
This action performs a one-shot image generation or editing operation directly based on the current prompt. It is typically used when the input description is simple, unambiguous, and requires no iterative refinement.
Prompt Enhancement with CoT.
This action refines the input prompt by enriching vague or underspecified user queries with additional contextual and descriptive details. Since text-to-image (T2I) models are highly sensitive to prompt wording [brade2023promptify, mo2024dynamic, liu2025one], more elaborate and specific prompts typically lead to higher-quality visual outputs. However, most T2I backbones are trained on simple captions or surface-level image descriptions, which limits their ability to reason over complex or compositional instructions. To address this limitation, we leverage the language reasoning capability of the model’s understanding module through Chain-of-Thought (CoT) prompting. By explicitly performing intermediate reasoning and elaboration, the agent transfers structured linguistic insights from the understanding domain to the generation process, thus producing more semantically aligned and visually coherent outputs.
Prompt Revision Based on the Generated/Edited Image.
This action is triggered when the generated or edited image indicates that the current prompt is suboptimal. In this case, the unified multimodal model is prompted to analyze the discrepancy between the visual output and the intended semantics, and to self-revise the prompt accordingly. This self-correction mechanism allows the agent to iteratively refine textual descriptions based on visual feedback, thereby enhancing both semantic alignment and generation quality over successive iterations.
Image Detail Refinement.
This action aims to correct minor imperfections in the generated or edited image when the input instruction is already satisfactory and the remaining issues stem from the generation module itself. It refines local visual details—such as textures, lighting, or small artifacts—without modifying the current prompt. By enhancing fine-grained fidelity while maintaining semantic consistency, this action improves the overall perceptual quality and realism of the generated results.
Best-of-N Sampling.
This action mitigates the inherent stochasticity of text-to-image (T2I) generation. Unlike language models that typically produce stable, low-entropy outputs through supervised fine-tuning (SFT) and reinforcement learning (RL) optimization, T2I models exhibit substantial variance across different samples [tian2025unigen, verdun2025soft, li2025reflect]. To reduce this randomness, the agent generates candidate images and employs the unified multimodal model to evaluate their visual and semantic alignment. The image with the highest alignment score is selected as the final output.
Termination (STOP).
This action signals the end of the iterative reasoning process when the agent determines that the current image and prompt have achieved satisfactory alignment. Instead of relying on a fixed number of steps, the agent autonomously decides to terminate based on its internal evaluation of visual quality.
| Types | Model | Cultural | Time | Space | Biology | Physics | Chemistry | Overall |
| SDv1.5 | 0.34 | 0.35 | 0.32 | 0.28 | 0.29 | 0.21 | 0.32 | |
| SDXL | 0.43 | 0.48 | 0.47 | 0.44 | 0.45 | 0.27 | 0.43 | |
| SD3.5-large | 0.44 | 0.50 | 0.58 | 0.44 | 0.52 | 0.31 | 0.46 | |
| PixArt-Alpha | 0.45 | 0.50 | 0.48 | 0.49 | 0.56 | 0.34 | 0.47 | |
| playground-v2.5 | 0.49 | 0.58 | 0.55 | 0.43 | 0.48 | 0.33 | 0.49 | |
| Gen | FLUX.1-dev | 0.48 | 0.58 | 0.62 | 0.42 | 0.51 | 0.35 | 0.50 |
| Janus | 0.16 | 0.26 | 0.35 | 0.28 | 0.30 | 0.14 | 0.23 | |
| VILA-U | 0.26 | 0.33 | 0.37 | 0.35 | 0.39 | 0.23 | 0.31 | |
| Show-o | 0.28 | 0.40 | 0.48 | 0.30 | 0.46 | 0.30 | 0.35 | |
| Emu3 | 0.34 | 0.45 | 0.48 | 0.41 | 0.45 | 0.27 | 0.39 | |
| MetaQuery-XL | 0.56 | 0.55 | 0.62 | 0.49 | 0.63 | 0.41 | 0.55 | |
| Unified | GPT-4o | 0.81 | 0.71 | 0.89 | 0.83 | 0.79 | 0.74 | 0.80 |
| Vanilla | 0.44 | 0.55 | 0.68 | 0.44 | 0.60 | 0.39 | 0.52 | |
| Bagel | ImAgent | 0.63 | 0.63 | 0.72 | 0.59 | 0.69 | 0.53 | 0.63 |
| Vanilla | 0.30 | 0.37 | 0.49 | 0.36 | 0.42 | 0.26 | 0.35 | |
| Janus-Pro-7B | ImAgent | 0.44 | 0.46 | 0.55 | 0.45 | 0.50 | 0.27 | 0.44 |
| Idiom | Textual | Entity | Scientific | Overall | |||||||
| Types | Model | Acc. | Qual. | Acc. | Qual. | Acc. | Qual. | Acc. | Qual. | Acc. | Qual. |
| FLUX.1-dev | 39.1 | 83.4 | 56.9 | 76.5 | 45.1 | 90.6 | 46.7 | 80.9 | 47.0 | 82.8 | |
| FLUX.1-schnell | 40.9 | 83.1 | 65.1 | 74.5 | 44.8 | 91.5 | 50.7 | 83.0 | 50.4 | 83.0 | |
| playground-v2.5 | 43.9 | 87.8 | 38.5 | 72.1 | 48.4 | 92.4 | 50.8 | 83.3 | 45.4 | 83.9 | |
| SD-3-Medium | 35.9 | 81.4 | 60.9 | 71.3 | 42.4 | 90.1 | 50.9 | 81.7 | 47.5 | 81.1 | |
| SD-3.5-Medium | 34.4 | 80.6 | 58.0 | 70.1 | 44.8 | 92.1 | 49.9 | 83.0 | 46.8 | 81.4 | |
| SD-3.5-Large | 35.6 | 85.3 | 62.2 | 75.4 | 46.6 | 92.6 | 52.9 | 84.5 | 49.3 | 84.4 | |
| Gen | gpt-image-1 | 75.7 | 94.5 | 86.9 | 97.6 | 77.5 | 96.6 | 74.7 | 94.3 | 78.7 | 95.8 |
| Emu3 | 33.1 | 82.9 | 33.7 | 68.7 | 33.8 | 85.2 | 40.1 | 77.0 | 35.2 | 78.5 | |
| Janus-Pro-7B | 25.5 | 78.0 | 37.2 | 70.9 | 38.5 | 87.6 | 44.9 | 77.8 | 36.5 | 78.6 | |
| show-o | 33.1 | 82.5 | 35.3 | 80.3 | 34.9 | 87.4 | 41.6 | 76.6 | 36.2 | 81.7 | |
| GoT | 29.7 | 76.4 | 30.6 | 70.7 | 31.0 | 86.2 | 36.8 | 76.3 | 32.0 | 77.4 | |
| Unified | Gemini-2.0 | 52.4 | 87.8 | 73.0 | 83.3 | 67.0 | 94.3 | 66.7 | 89.3 | 64.8 | 88.7 |
| Vanilla | 30.2 | 85.7 | 36.6 | 68.4 | 45.0 | 94.7 | 54.4 | 87.5 | 41.6 | 84.1 | |
| Bagel | ImAgent | 37.7 | 90.0 | 54.2 | 79.1 | 52.6 | 96.6 | 61.2 | 90.3 | 51.4 | 89.0 |
| Vanilla | 25.5 | 78.0 | 37.2 | 70.9 | 38.5 | 87.6 | 44.9 | 77.8 | 36.5 | 78.6 | |
| Janus-Pro-7B | ImAgent | 27.9 | 86.0 | 35.3 | 68.8 | 40.7 | 89.7 | 51.2 | 84.2 | 38.8 | 82.2 |
4 Experiments
4.1 Experimental Setting
Models.
We build our ImAgent on two of the most recent and powerful unified models, Bagel [deng2025emerging] and Janus-Pro-7B [chen2025janus]. Both models possess versatile multimodal capabilities, including image understanding, image generation, which collectively form the foundation of our framework. For image generation, we leverage both Bagel and Janus-Pro-7B. However, since Janus-Pro-7B does not support image editing, we employ Bagel exclusively for the image editing tasks.
Benchmarks.
We evaluate ImAgent on multiple benchmarks, including 3 benchmarks on image generation and 4 benchmarks for image editing tasks. For image generation, we use R2I-Bench [chen2025r2i], which is designed to rigorously assess reasoning-driven T2I generation; WISE [niu2025wise], which covers six categories of image generation scenarios; and T2I-ReasonBench [sun2025t2i], which assesses the models’ reasoning ability in generative tasks. For image editing, we adopt GEdit-Bench [liu2504step1x], which contains both Chinese and English instruction-based editing tasks; RISEBench [zhao2025envisioning], which focuses on reasoning-informed visual editing across diverse reasoning types; KRIS-Bench [wu2025kris], which evaluates reasoning capabilities over factual, conceptual, and procedural knowledge; and ImgEdit-Bench [ye2025imgedit], which is used to evaluate image editing performance in terms of instruction adherence, editing quality, and detail preservation.
Baselines.
For image generation, we employ SDv1.5 [rombach2022high], SD3-Medium [rombach2022high], SDXL [podell2023sdxl], SD3.5-Medium [esser2024scaling], SD3.5-Large [esser2024scaling], PixArt-Alpha [chen2024pixart], FLUX.1-Dev [blackforest2024flux], Sana-1.5 [xie2025sana], Lumina-T2I [qin2025lumina], LLM4GENSD1.5 [liu2025llm4gen], ELLASD1.5 [hu2024ella], LlamaGen [sun2024autoregressive], DALL-E-3 [ma2025learning], gpt-image-1 [hurst2024gpt], Omnigen [xiao2025omnigen], FLUX.1-schnell [flux2024], andPlayground-v2.5 [li2024playground], as generation-based models. We also include Janus [wu2025janus], VILA-U [wu2024vila], Show-o [xie2024show], Janus-Pro-7B [chen2025janus], Emu3 [wang2024emu3], Lumina-Image 2.0 [qin2025lumina], show-o [xie2024show], GoT [fang2025got] and MetaQuery-XL [pan2025transfer] as unified multimodal models.
| Types | Model | Temporal | Causal | Spatial | Logical | Overall |
| Gemini-2.0-Flash-pre | 10.6 | 13.3 | 11.0 | 2.3 | 9.4 | |
| Gemini-2.0 | 8.2 | 15.5 | 23.0 | 4.7 | 13.3 | |
| Private | GPT-4o | 34.1 | 32.2 | 37.0 | 10.6 | 28.9 |
| EMU2 | 1.2 | 1.1 | 0.0 | 0.0 | 0.5 | |
| OmniGen | 1.2 | 1.0 | 0.0 | 1.2 | 0.8 | |
| Step1X-Edit | 0.0 | 2.2 | 2.0 | 3.5 | 1.9 | |
| Qwen-Image-Edit | 4.7 | 10.0 | 17.0 | 2.4 | 8.9 | |
| FLUX.1-Kontext-Dev | 2.3 | 5.5 | 13.0 | 1.2 | 5.8 | |
| Ovis-U1 | 1.2 | 3.3 | 4.0 | 2.4 | 2.8 | |
| Open | Seedream-4.0 | 12.9 | 12.2 | 11.0 | 7.1 | 10.8 |
| Vanilla | 2.4 | 5.6 | 14.0 | 1.2 | 6.1 | |
| Bagel | ImAgent | 17.6 | 15.6 | 16.0 | 2.4 | 13.1 |
For image editing, we evaluate both private and open baselines. The private baselines include Gemini-2.0 [google2025_gemini2_image], Doubao [shi2411seededit], GPT-4o [openai2025gpt4o_image], and Gemini-2.0-Flash-pre [team2023gemini]. The open baselines include Step1X-Edit [liu2025step1x], Instruct-Pix2Pix [brooks2023instructpix2pix], MagicBrush [zhang2023magicbrush], AnyEdit [yu2025anyedit], OmniGen [xiao2025omnigen], EMU2 [sun2024generative], Qwen-Image-Edit [wu2025qwenimagetechnicalreport], FLUX.1-Kontext-Dev [labs2025flux1kontextflowmatching], Ovis-U1 [wang2025ovisu1], and Seedream-4.0 [seedream2509seedream].
4.2 Quantitative Results
Image Generation.
As shown in Table 2, ImAgent achieves outstanding performance on the WISE benchmark. Compared to their respective backbone models, ImAgent built upon Bagel and Janus-Pro-7B consistently surpasses the vanilla counterparts, with performance improvements of 21.2% and 25.7%, respectively, demonstrating the effectiveness of our approach in enhancing image generation quality through coordinated reasoning and refinement. Notably, the vanilla Janus-Pro-7B underperforms compared to SDXL and EMU3 and performs on par with Show-o. In contrast, ImAgent constructed on Janus-Pro-7B surpasses all of them, highlighting its strong test-time scaling capability and generalization potential. Furthermore, on the reasoning benchmarks, ImAgent consistently delivers strong results. As shown in Tables 1 and 3, ImAgent outperforms the vanilla models across both backbone architectures. For instance, ImAgent achieves improvements of 14.8% and 7.5% across both backbones on R2I-Bench, respectively. Moreover, ImAgent built upon Bagel surpasses the commercial model Gemini-2.0 in overall quality on the T2I-ReasonBench, further demonstrating its effectiveness in reasoning-based image generation.
Image Editing.
As shown in Table 4, vanilla Bagel achieves only 6.1 on RISEBench, underperforming compared to Qwen-Image-Edit and Seedream-4.0. In contrast, ImAgent achieves a 114.8% improvement over vanilla and surpasses both Qwen-Image-Edit and Seedream-4.0. Remarkably, ImAgent performs on par with Gemini-2.0, demonstrating its effectiveness and strong test-time scaling capability. This indicates that open-source models enhanced with ImAgent can achieve competitive performance with commercial models.
Furthermore, as shown in Figure 4, ImAgent also achieves excellent performance on KRISBench, attaining an overall score of 67.13, surpassing vanilla Bagel which scores 63.16. ImAgent performs well on ImgEdit-Bench (shown in Figure 3) as well, with vanilla achieving 2.89 and underperforming compared to Step1X-Edit, while ImAgent reaches 3.15, outperforming Step1X-Edit and achieving the best performance among open-source models. This further demonstrates the effectiveness of ImAgent in test-time scaling.
Moreover, ImAgent outperforms vanilla Bagel across both English and Chinese settings in GEdit-Bench, achieving improvements of 5.5% and 5.2%, respectively. Detailed results are provided in the Appendix.
4.3 Qualitative Analysis
Image Refinement Flow.
As shown in Figure 1, we present several qualitative examples, including two for image generation and two for image editing, each executed for five steps. From an overall perspective, we can observe that the image quality improves progressively with each step, demonstrating that ImAgent refines the image gradually.
Specifically, consider a case where the user aims to generate an image based on the prompt “moldy bread.” In the first step, ImAgent performs the Naive Generation action and produces an image containing fresh bread, which clearly deviates from the prompt. The policy controller then selects the Prompt Refinement action to improve the prompt using both the generated image and the user’s original input. Although the updated image shows slight improvement, it remains unsatisfactory. To address this, the controller chooses the Image Detail Refinement action, leading to the appearance of mold on the bread. Subsequently, ImAgent applies the Prompt Enhancement action to further strengthen the textual guidance, resulting in an image that better aligns with the user’s intent. Finally, recognizing that the prompt is now sufficiently clear but the visual result could still be improved, ImAgent executes the Best-of-N Sampling action to generate multiple candidates and select the most aligned one. The final output successfully captures the desired “moldy bread” concept.
Detailed Policy Controller Decision.
We present a detailed case study of the policy controller’s output to illustrate how it makes action decisions. As shown in Figure 6, which visualizes the decision-making process for the case discussed earlier, the policy controller selects the promptrefinement action to enrich the textual prompt and improve semantic clarity for subsequent generation steps. We present more examples in the Appendix.
4.4 Ablation Studies
Comparison with Self-CoT.
Since the vanilla Bagel model already incorporates self-CoT capabilities, we further compare ImAgent with self-CoT to evaluate the effectiveness of our approach. For the image editing task, as shown in Figure 5(a), ImAgent significantly outperforms both the vanilla and self-CoT variants on the Gedit-Bench across both English and Chinese settings. Furthermore, ImAgent also surpasses self-CoT in the image generation task. As illustrated in Figures 5(b) and (c), under the Janus-Pro-7B backbone, the performance of self-CoT notably drops compared to the vanilla model, whereas ImAgent consistently achieves superior results. Similarly, under the Bagel backbone, ImAgent maintains clear advantages over both the vanilla and self-CoT models. These findings suggest that the improvements brought by ImAgent stem not only from self-CoT–like reasoning refinement, but also from the different actions coordination enabled by the strong policy controller.
Effectiveness of the Whole ImAgent.
To evaluate the effectiveness of ImAgent in coordinately selecting and executing appropriate actions, we compare its performance with that of single-turn actions and a random action selection policy, where the for random selection is set the same as in ImAgent. As shown in Table 5, experiments are conducted on R2I-Bench [chen2025r2i]. All single-turn actions perform well compared to the naive generation action, which serves as the lower bound (0.54). Moreover, the random action selection policy achieves a score of 0.59, higher than all single-turn actions, indicating that naively constructing an agent is beneficial. Finally, ImAgent outperforms the random policy with a score of 0.62, demonstrating that it effectively and efficiently boosts test-time scaling in image generation.
| Method | Random | ImAgent | |||||
| Score | 0.54 | 0.58 | 0.57 | 0.56 | 0.56 | 0.59 | 0.62 |
5 Conclusion
In this paper, we presented ImAgent, a training-free unified multimodal agent designed to enhance image generation through efficient test-time scaling. By integrating multiple generation actions within a single framework and coordinating them via a policy controller, ImAgent dynamically selects and executes the most appropriate action for a given case. Extensive experiments on image generation and editing benchmarks demonstrate that ImAgent consistently improves over the backbone models and surpasses strong baseline methods, including scenarios where the backbone fails. These results highlight the potential of unified multimodal agents for adaptive, efficient, and high-quality image generation without relying on additional external models.
Supplementary Material
Appendix A Algorithm for Image Editing
We present the detailed algorithm (shown in Algorithm 2) for image editing.
Appendix B Detailed Action Description
Naive Generation/Editing.
This action performs a direct, one-shot invocation of the underlying text-to-image or image-editing backbone without any intermediate reasoning or prompt restructuring. Given the current prompt–image pair , the backbone generator produces the next image deterministically as:
This action is effective when the user intent is already explicit, the visual goal is straightforward, or iterative refinement offers limited benefit.
Prompt Enhancement with CoT.
This action enriches the input prompt using the model’s multimodal reasoning capability. Unlike naive generation, it introduces an intermediate reasoning stage, where the model generates a chain-of-thought (CoT) enhanced prompt that clarifies ambiguous instructions, decomposes complex compositions, and fills missing semantic details:
The intuition is that while T2I models excel at visual synthesis, they often struggle with compositional or underspecified instructions due to their training bias toward short captions. By transferring the model’s linguistic reasoning ability to the generative pathway, CoT enhancement produces prompts that are more explicit, structured, and visually actionable. Finally, the updated prompt is fed into the generator:
This action is particularly beneficial when the input contains complex spatial relationships, multi-object instructions, stylistic requirements, or vague conceptual queries. By exposing the implicit reasoning steps, the model produces outputs that are more semantically aligned, compositionally consistent, and visually faithful to user intent.
| Method | Factual | Conceptual | Procedural | Overall | |||
| AP | SP | SS | NS | LP | ID | ||
| Vanilla | 73.27 | 67.17 | 61.10 | 62.36 | 54.38 | 64.77 | 63.16 |
| ImAgent | 74.39 | 71.15 | 69.80 | 66.03 | 54.83 | 68.78 | 67.13 |
| Types | Models | Add | Adjust | Extract | Replace | Remove | Background | Style | Hybrid | Action | Overall |
| Private | GPT-4o-Image | 4.61 | 4.33 | 2.90 | 4.35 | 3.66 | 4.57 | 4.93 | 3.96 | 4.89 | 4.20 |
| Open | MagicBrush | 2.84 | 1.58 | 1.51 | 1.97 | 1.58 | 1.75 | 2.38 | 1.62 | 1.22 | 1.83 |
| Instruct-P2P | 2.45 | 1.83 | 1.44 | 2.01 | 1.50 | 1.44 | 3.55 | 1.20 | 1.46 | 1.88 | |
| AnyEdit | 3.18 | 2.95 | 1.88 | 2.47 | 2.23 | 2.24 | 2.85 | 1.56 | 2.65 | 2.45 | |
| UltraEdit | 3.44 | 2.81 | 2.13 | 2.96 | 1.45 | 2.83 | 3.76 | 1.91 | 2.98 | 2.70 | |
| Step1X-Edit | 3.88 | 3.14 | 1.76 | 3.40 | 2.41 | 3.16 | 4.63 | 2.64 | 2.52 | 3.06 | |
| UniWorld-V1 | 3.82 | 3.64 | 2.27 | 3.47 | 3.24 | 2.99 | 4.21 | 2.96 | 2.74 | 3.26 | |
| Bagel | Vanilla | 3.89 | 3.68 | 1.33 | 3.78 | 3.07 | 2.07 | 4.30 | 2.47 | 4.33 | 2.89 |
| ImAgent | 3.20 | 3.89 | 1.77 | 3.13 | 2.71 | 3.50 | 4.21 | 2.89 | 4.43 | 3.15 |
Prompt Revision Based on the Generated/Edited Image.
This action is invoked when the current visual output reveals a mismatch between the generated image and the intended semantics. In such cases, the unified multimodal model performs discrepancy-aware prompt revision. Specifically, the model first extracts a semantic summary from the generated image:
and evaluates its alignment with the intended description encoded in the prompt:
where captures the semantic gap (e.g., missing attributes, incorrect object relations). The revised prompt is then formulated as:
This action enables the agent to iteratively correct textual instructions based on visual feedback. This self-correcting loop strengthens consistency between linguistic intent and visual realization, leading to progressively refined generations across steps.
Image Detail Refinement.
This action focuses on enhancing the fidelity of local image details while keeping the semantic content unchanged. When the prompt is already accurate but the model output exhibits minor imperfections—such as texture noise, inconsistent lighting, or small structural artifacts—the agent performs detail-aware refinement. Formally, given the current image , a refinement operator is applied:
where leverages the prompt only as a semantic constraint to prevent drift. This action improves perceptual quality without altering scene composition, making it especially useful in polishing near-final results.
Best-of-N Sampling.
This action addresses the high stochasticity inherent in T2I generation, where multiple plausible images may arise from the same prompt. To reduce variance, the agent produces a set of candidate images:
and the unified multimodal model assigns each candidate an alignment score:
reflecting semantic consistency and visual quality. The final selected output is:
This sampling strategy stabilizes generation quality and significantly mitigates randomness caused by diffusion or autoregressive sampling.
Appendix C Detailed Experimental Results
Appendix D Detailed Policy Controller Decision
We present a case study that illustrates the step-by-step decisions made by the policy controller. While Section 4.3 provides a partial walkthrough of how ImAgent operates on a given example, here we provide the complete decision trajectory, including all intermediate actions, observations, and updates performed by the agent throughout the entire process.
For the first step, ImAgent decides to invoke the Naive Generation action, here is the detailed reason:
For the second step, ImAgent decides to invoke the Prompt Refinement:
For the third step, ImAgent decides to invoke the Image Detail Refinement:
For the forth step, ImAgent decides to invoke the Prompt Enhancement:
For the fifth step, ImAgent decides to invoke the Best-of-N Sampling:
| Types | Models | EN Setting | CN Setting | ||||
| G_SC | G_PQ | G_O | G_SC | G_PQ | G_O | ||
| Private | Gemini-2.0 | 6.73 | 6.61 | 6.32 | 5.43 | 6.77 | 5.36 |
| Doubao | 6.92 | 7.19 | 6.75 | 6.98 | 7.27 | 6.77 | |
| GPT-4o | 7.85 | 7.62 | 7.53 | 7.67 | 7.56 | 7.30 | |
| Open | Instruct-Pix2Pix | 3.58 | 5.50 | 3.68 | - | - | - |
| MagicBrush | 4.68 | 5.66 | 4.52 | - | - | - | |
| AnyEdit | 3.18 | 5.82 | 3.21 | - | - | - | |
| OmniGen | 5.96 | 5.89 | 5.06 | - | - | - | |
| Step1X-Edit | 7.09 | 6.76 | 6.70 | 7.20 | 6.87 | 6.86 | |
| Bagel | Vanilla | 7.36 | 6.83 | 6.52 | 7.34 | 6.85 | 6.50 |
| ImAgent | 7.85 | 6.61 | 6.88 | 7.92 | 6.53 | 6.84 | |

### v3

@@
 1 Introduction
-Text-to-Image (T2I) models [rombach2022high, ramesh2022hierarchical, podell2023sdxl, sun2024autoregressive, tian2024visual, qin2025lumina] have made remarkable progress in generating visually realistic and semantically coherent images from natural language descriptions.
-Despite these advancements, the quality of generated results often exhibits randomness and inconsistency with the given prompts [xu2024prompt, lee2024direct].
+Text-to-Image (T2I) models [36, 35, 33, 44, 47, 34] have made remarkable progress in generating visually realistic and semantically coherent images from natural language descriptions.
+Despite these advancements, the quality of generated results often exhibits randomness and inconsistency with the given prompts [62, 18].
 This limitation primarily stems from the strong dependency of existing T2I models on the clarity and specificity of textual inputs.
 When the prompt is vague or underspecified, the models tend to overlook key semantic elements, leading to images that deviate from the intended meaning and fail to fully capture the user’s intent.
-To mitigate this problem, prior research has proposed various strategies, such as prompt rewriting [brade2023promptify, mo2024dynamic, liu2025one], best-of-N sampling [li2025reflect, singhal2025general], classifier-free guidance [ho2022classifier, sauer2023stylegan, shen2024rethinking], and self-revision or iterative refinement [singh2023divide, li2024g, jeon2025iterative].
+To mitigate this problem, prior research has proposed various strategies, such as prompt rewriting [1, 29, 25], best-of-N sampling [21, 42], classifier-free guidance [11, 37, 39], and self-revision or iterative refinement [41, 19, 14].
 These approaches aim to either clarify the textual prompt or reduce the model’s inherent randomness rather than finetuning the models, reflecting the underlying principle of test-time scaling—improving generation quality through additional inference-time computation.
@@
 This design enables more efficient test-time scaling.
-Thanks to recent advances in unified multimodal models [deng2025emerging, wu2025janus, wu2025janus, xie2024show, qin2025lumina], which integrate text generation, image generation, and visual understanding within a single framework, this assumption becomes attainable.
+Thanks to recent advances in unified multimodal models [6, 55, 55, 59, 34], which integrate text generation, image generation, and visual understanding within a single framework, this assumption becomes attainable.
 Building upon this foundation, we introduce ImAgent, a training-free unified multimodal agent designed to perform efficient test-time scaling for image generation.
@@
 Text-to-Image (T2I) models have demonstrated remarkable ability in synthesizing visually realistic and semantically coherent images from natural language descriptions.
-The evolution of T2I models has followed several major paradigms, including diffusion-based models [podell2023sdxl, rombach2022high, ramesh2022hierarchical, esser2024scaling, gao2024lumina, xie2025sana], autoregressive models [sun2024autoregressive, zhang2024var, tian2024visual, wang2024emu3, zheng2025parallel, xiong2025llava], and more recently, unified multimodal models [xiao2025omnigen, xie2024show, zhou2024transfusion, tong2025metamorph, deng2025emerging, zheng2025learning, luo2024stable].
+The evolution of T2I models has followed several major paradigms, including diffusion-based models [33, 36, 35, 7, 9, 58], autoregressive models [44, 69, 47, 52, 73, 61, 60], and more recently, unified multimodal models [57, 59, 74, 49, 6, 72, 26].
 These approaches have significantly advanced image fidelity, diversity, and text-image alignment.
@@
 Unified multimodal models refer to models that jointly possess multimodal understanding and generation capabilities within a single framework.
-Unlike traditional pipelines that rely on separate models for text or image understanding, image generation, and image editing, these models integrate vision and language processing into a shared architecture, enabling seamless information exchange across modalities [xiao2025omnigen, zhou2024transfusion, tong2025metamorph, wu2025janus, wu2025janus].
+Unlike traditional pipelines that rely on separate models for text or image understanding, image generation, and image editing, these models integrate vision and language processing into a shared architecture, enabling seamless information exchange across modalities [57, 74, 49, 55, 55].
 This convergence of understanding and generation not only improves efficiency and coherence but also paves the way for building more general-purpose, human-like agents capable of reasoning, creating, and self-correcting.
@@
 To alleviate these issues, various optimization strategies have been proposed.
-Best-of-N sampling [tian2025unigen, verdun2025soft, li2025reflect] generates multiple candidates and selects the best one, reducing the influence of stochastic variations in model sampling.
-Prompt rewriting [tian2025unigen, wang2024promptcharm, yasunaga2022retrieval] reformulates or enriches the input text to provide more explicit guidance for the generator, improving semantic alignment and visual coherence.
-classifier-free guidance [ho2022classifier, sauer2023stylegan, shen2024rethinking] controls the conditioning strength to balance fidelity and diversity, and iterative self-refinement [yang2023idea2img, madaan2023self, singh2023divide, li2024g, jeon2025iterative] enables models to assess and revise their outputs through multiple rounds of generation.
+Best-of-N sampling [48, 50, 21] generates multiple candidates and selects the best one, reducing the influence of stochastic variations in model sampling.
+Prompt rewriting [48, 53, 64] reformulates or enriches the input text to provide more explicit guidance for the generator, improving semantic alignment and visual coherence.
+classifier-free guidance [11, 37, 39] controls the conditioning strength to balance fidelity and diversity, and iterative self-refinement [63, 28, 41, 19, 14] enables models to assess and revise their outputs through multiple rounds of generation.
 Although these techniques significantly enhance performance, they typically operate as independent modules with limited interaction between reasoning, generation, and evaluation—highlighting the need for unified multimodal frameworks that integrate these capabilities into a cohesive pipeline.
@@
 We construct ImAgent upon unified multimodal models, which inherently possess both understanding and generation capabilities.
-Motivated by the observation that multimodal understanding is generally more reliable and easier to achieve than multimodal generation [zhang2025unified, zhang2025unified1], ImAgent leverages the model’s strong understanding ability to adaptively select the optimal action for a given case that enhances the overall generation quality.
+Motivated by the observation that multimodal understanding is generally more reliable and easier to achieve than multimodal generation [70, 67], ImAgent leverages the model’s strong understanding ability to adaptively select the optimal action for a given case that enhances the overall generation quality.
 This design enables ImAgent to fully exploit the potential of a single unified multimodal model without any additional training or external modules, thereby achieving efficient test-time scaling.
@@
 Unlike conventional agent-based frameworks where observations directly influence state updates, in ImAgent the observation history is primarily used for action selection, while prompt and image updates are determined by the selected generation actions.
-This iterative process continues until the agent selects the STOP action, which indicates that the generated result is satisfactory, or until the maximum number of steps (set to 5 by default) is reached.
+This iterative process continues until the agent selects the STOP action, which indicates that the generated result is satisfactory, or until the maximum number of steps is reached.
 Notably, both the policy controller and the action execution are implemented within a single unified multimodal model.
 The detailed algorithm for image generation is presented in Algorithm 1, while the algorithm for image editing follows a similar procedure and is provided in the Appendix.
+3.2 Action Space
+In this section, we introduce the action space of ImAgent.
+The detailed definitions and usage of each action are provided in the Appendix 7.
+The detailed prompts for each action are provided in the Appendix 9.
+Naive Generation/Editing.
+This action performs a one-shot image generation or editing operation directly based on the current prompt.
+It is typically used when the input description is simple, unambiguous, and requires no iterative refinement.
+Prompt Enhancement with CoT.
+This action refines the input prompt by enriching vague or underspecified user queries with additional contextual and descriptive details.
+Since text-to-image (T2I) models are highly sensitive to prompt wording [1, 29, 25], more elaborate and specific prompts typically lead to higher-quality visual outputs.
+However, most T2I backbones are trained on simple captions or surface-level image descriptions, which limits their ability to reason over complex or compositional instructions.
+To address this limitation, we leverage the language reasoning capability of the model’s understanding module through Chain-of-Thought (CoT) prompting.
+By explicitly performing intermediate reasoning and elaboration, the agent transfers structured linguistic insights from the understanding domain to the generation process, thus producing more semantically aligned and visually coherent outputs.
+Prompt Revision Based on the Generated/Edited Image.
+This action is triggered when the generated or edited image indicates that the current prompt is suboptimal.
+In this case, the unified multimodal model is prompted to analyze the discrepancy between the visual output and the intended semantics, and to self-revise the prompt accordingly.
+This self-correction mechanism allows the agent to iteratively refine textual descriptions based on visual feedback, thereby enhancing both semantic alignment and generation quality over successive iterations.
+Image Detail Refinement.
+This action targets minor visual imperfections when the input instruction is already satisfactory and the remaining issues mainly originate from the generation module.
+It focuses on refining local visual details such as textures, lighting, and small artifacts.
+Specifically, the policy controller automatically generates a local editing prompt conditioned on the current image and the original user prompt, which is used internally to guide fine-grained image editing.
+Note that this local editing prompt is different from the user-provided prompt in image editing tasks.
+By enhancing fine-grained visual fidelity while preserving semantic consistency, this action improves the perceptual quality and realism of the generated results.
+Best-of-N Sampling.
+This action mitigates the inherent stochasticity of T2I generation.
+Unlike language models that typically produce stable, low-entropy outputs through supervised fine-tuning (SFT) and reinforcement learning (RL) optimization, T2I models exhibit substantial variance across different samples [48, 50, 21].
+To reduce this randomness, this action generates candidate images and then evaluate their visual and semantic alignment based on the given prompt.
+The image with the highest alignment score is selected as the final output.
+STOP.
+This action signals the end of the iterative reasoning process when the agent determines that the current image and prompt have achieved satisfactory alignment.
+Instead of relying on a fixed number of steps, the agent autonomously decides to terminate based on its internal evaluation of visual quality.
+3.3 Policy Controller
+Decision Prompt.
+We construct a structured decision prompt for the policy controller, which takes the current prompt, the generated image, and the observation history as inputs, and outputs a discrete action from the predefined action space .
+As shown in Figure 3, the decision prompt includes the current generation prompt and the observation history (including action and prompt histories), and explicitly instructs the model to select the most suitable next action from the predefined action space.
+Action Parsing.
+[…]
+To mitigate hallucinated or invalid action outputs, we adopt a fallback strategy: if the parsed action is not in the predefined action set, we default to the Naive Generation/Editing action.
+This design ensures stable execution even when the policy controller produces malformed or out-of-distribution outputs.
+We provide a comprehensive robustness analysis of ImAgent in Section 4.3.
+4 Experiments
+4.1 Experimental Setting
 | Types | Model | Comm. | Comp. | Logical | Numerical | Causal | Overall |
@@
 | Unified | Lumina-Image 2.0 | 0.49 | 0.65 | 0.56 | 0.43 | 0.40 | 0.52 |
-| Vanilla | 0.47 | 0.65 | 0.53 | 0.68 | 0.40 | 0.54 | |
-| Bagel | ImAgent | 0.58 | 0.68 | 0.62 | 0.71 | 0.53 | 0.62 |
-| Vanilla | 0.45 | 0.60 | 0.46 | 0.46 | 0.36 | 0.47 | |
-| Janus-Pro-7B | ImAgent | 0.47 | 0.58 | 0.52 | 0.54 | 0.44 | 0.51 |
-3.2 Action Space
-In this section, we introduce the action space of ImAgent. The detailed definitions and usage of each action are provided in the Appendix.
-Naive Generation/Editing.
-This action performs a one-shot image generation or editing operation directly based on the current prompt.
-It is typically used when the input description is simple, unambiguous, and requires no iterative refinement.
-Prompt Enhancement with CoT.
-This action refines the input prompt by enriching vague or underspecified user queries with additional contextual and descriptive details.
-Since text-to-image (T2I) models are highly sensitive to prompt wording [brade2023promptify, mo2024dynamic, liu2025one], more elaborate and specific prompts typically lead to higher-quality visual outputs.
-However, most T2I backbones are trained on simple captions or surface-level image descriptions, which limits their ability to reason over complex or compositional instructions.
-To address this limitation, we leverage the language reasoning capability of the model’s understanding module through Chain-of-Thought (CoT) prompting.
-By explicitly performing intermediate reasoning and elaboration, the agent transfers structured linguistic insights from the understanding domain to the generation process, thus producing more semantically aligned and visually coherent outputs.
-Prompt Revision Based on the Generated/Edited Image.
-This action is triggered when the generated or edited image indicates that the current prompt is suboptimal.
-In this case, the unified multimodal model is prompted to analyze the discrepancy between the visual output and the intended semantics, and to self-revise the prompt accordingly.
-This self-correction mechanism allows the agent to iteratively refine textual descriptions based on visual feedback, thereby enhancing both semantic alignment and generation quality over successive iterations.
-Image Detail Refinement.
-This action aims to correct minor imperfections in the generated or edited image when the input instruction is already satisfactory and the remaining issues stem from the generation module itself.
-It refines local visual details—such as textures, lighting, or small artifacts—without modifying the current prompt.
-By enhancing fine-grained fidelity while maintaining semantic consistency, this action improves the overall perceptual quality and realism of the generated results.
-Best-of-N Sampling.
-This action mitigates the inherent stochasticity of text-to-image (T2I) generation.
-Unlike language models that typically produce stable, low-entropy outputs through supervised fine-tuning (SFT) and reinforcement learning (RL) optimization, T2I models exhibit substantial variance across different samples [tian2025unigen, verdun2025soft, li2025reflect].
-To reduce this randomness, the agent generates candidate images and employs the unified multimodal model to evaluate their visual and semantic alignment.
-The image with the highest alignment score is selected as the final output.
-Termination (STOP).
-This action signals the end of the iterative reasoning process when the agent determines that the current image and prompt have achieved satisfactory alignment.
-Instead of relying on a fixed number of steps, the agent autonomously decides to terminate based on its internal evaluation of visual quality.
+| Bagel | Vanilla | 0.47 | 0.65 | 0.53 | 0.68 | 0.40 | 0.54 |
+| Self-Rewrite | 0.54 | 0.67 | 0.54 | 0.69 | 0.47 | 0.58 | |
+| ImAgent | 0.58 | 0.68 | 0.62 | 0.71 | 0.53 | 0.62 | |
+| Janus-Pro-7B | Vanilla | 0.45 | 0.60 | 0.46 | 0.46 | 0.36 | 0.47 |
+| Self-Rewrite | 0.40 | 0.48 | 0.35 | 0.39 | 0.28 | 0.38 | |
+| ImAgent | 0.47 | 0.58 | 0.52 | 0.54 | 0.44 | 0.51 |
+| Types | Model | Temporal | Causal | Spatial | Logical | Overall |
+| Private | Gemini-2.0-Flash-pre | 10.6 | 13.3 | 11.0 | 2.3 | 9.4 |
+| Gemini-2.0 | 8.2 | 15.5 | 23.0 | 4.7 | 13.3 | |
+| GPT-4o | 34.1 | 32.2 | 37.0 | 10.6 | 28.9 | |
+| Open | EMU2 | 1.2 | 1.1 | 0.0 | 0.0 | 0.5 |
+| OmniGen | 1.2 | 1.0 | 0.0 | 1.2 | 0.8 | |
+| Step1X-Edit | 0.0 | 2.2 | 2.0 | 3.5 | 1.9 | |
+| Qwen-Image-Edit | 4.7 | 10.0 | 17.0 | 2.4 | 8.9 | |
+| FLUX.1-Kontext-Dev | 2.3 | 5.5 | 13.0 | 1.2 | 5.8 | |
+| Ovis-U1 | 1.2 | 3.3 | 4.0 | 2.4 | 2.8 | |
+| Seedream-4.0 | 12.9 | 12.2 | 11.0 | 7.1 | 10.8 | |
+| Bagel | Vanilla | 2.4 | 5.6 | 14.0 | 1.2 | 6.1 |
+| Self-Rewrite | 5.9 | 17.8 | 21.0 | 1.2 | 11.9 | |
+| ImAgent | 17.6 | 15.6 | 16.0 | 2.4 | 13.1 |
+Agentic Details.
+The maximum number of steps is set to 5 by default, meaning that the agent will automatically select the STOP action if this limit is reached. For the Best-of-N Sampling action, we set by default.
+Backbone Models.
+We build our ImAgent on two of the most recent and powerful unified models, Bagel [6] and Janus-Pro-7B [5].
+Both models possess versatile multimodal capabilities, including image understanding, image generation, which collectively form the foundation of our framework.
+For image generation, we leverage both Bagel and Janus-Pro-7B.
+However, since Janus-Pro-7B does not support image editing, we employ Bagel exclusively for the image editing tasks.
+Benchmarks.
+We evaluate ImAgent on multiple benchmarks, including 3 benchmarks on image generation and 3 benchmarks for image editing tasks.
+For image generation, we use R2I-Bench [4], which is designed to rigorously assess reasoning-driven T2I generation; WISE [30], which covers six categories of image generation scenarios; and T2I-ReasonBench [43], which assesses the models’ reasoning ability in generative tasks.
+For image editing, we adopt GEdit-Bench [23], which contains both Chinese and English instruction-based editing tasks; RISEBench [71], which focuses on reasoning-informed visual editing across diverse reasoning types; and ImgEdit-Bench [65], which is used to evaluate image editing performance in terms of instruction adherence, editing quality, and detail preservation.
+Baselines.
+For image generation, we employ SDv1.5 [36], SD3-Medium [36], SDXL [33], SD3.5-Medium [7], SD3.5-Large [7], PixArt-Alpha [3], FLUX.1-Dev [16], Sana-1.5 [58], Lumina-T2I [34], LLM4GENSD1.5 [22], ELLASD1.5 [12], LlamaGen [44], DALL-E-3 [27], gpt-image-1 [13], Omnigen [57], FLUX.1-schnell [17], andPlayground-v2.5 [20], as generation-based models.
+We also include Janus [55], VILA-U [56], Show-o [59], Janus-Pro-7B [5], Emu3 [52], Lumina-Image 2.0 [34], show-o [59], GoT [8] and MetaQuery-XL [32] as unified multimodal models.
+For image editing, we evaluate both private and open baselines.
+The private baselines include Gemini-2.0 [10], Doubao [40], GPT-4o [31], and Gemini-2.0-Flash-pre [46].
+The open baselines include Step1X-Edit [24], Instruct-Pix2Pix [2], MagicBrush [68], AnyEdit [66], OmniGen [57], EMU2 [45], Qwen-Image-Edit [54], FLUX.1-Kontext-Dev [15], Ovis-U1 [51], and Seedream-4.0 [38].
+Meanwhile, we also compare ImAgent with the backbone unified model and a prompt self-rewrite baseline implemented using the same backbone model.
 | Types | Model | Cultural | Time | Space | Biology | Physics | Chemistry | Overall |
-| SDv1.5 | 0.34 | 0.35 | 0.32 | 0.28 | 0.29 | 0.21 | 0.32 | |
+| Gen | SDv1.5 | 0.34 | 0.35 | 0.32 | 0.28 | 0.29 | 0.21 | 0.32 |
 | SDXL | 0.43 | 0.48 | 0.47 | 0.44 | 0.45 | 0.27 | 0.43 | |
@@
 | playground-v2.5 | 0.49 | 0.58 | 0.55 | 0.43 | 0.48 | 0.33 | 0.49 | |
-| Gen | FLUX.1-dev | 0.48 | 0.58 | 0.62 | 0.42 | 0.51 | 0.35 | 0.50 |
-| Janus | 0.16 | 0.26 | 0.35 | 0.28 | 0.30 | 0.14 | 0.23 | |
+| FLUX.1-dev | 0.48 | 0.58 | 0.62 | 0.42 | 0.51 | 0.35 | 0.50 | |
+| Unified | Janus | 0.16 | 0.26 | 0.35 | 0.28 | 0.30 | 0.14 | 0.23 |
 | VILA-U | 0.26 | 0.33 | 0.37 | 0.35 | 0.39 | 0.23 | 0.31 | |
@@
 | MetaQuery-XL | 0.56 | 0.55 | 0.62 | 0.49 | 0.63 | 0.41 | 0.55 | |
-| Unified | GPT-4o | 0.81 | 0.71 | 0.89 | 0.83 | 0.79 | 0.74 | 0.80 |
-| Vanilla | 0.44 | 0.55 | 0.68 | 0.44 | 0.60 | 0.39 | 0.52 | |
-| Bagel | ImAgent | 0.63 | 0.63 | 0.72 | 0.59 | 0.69 | 0.53 | 0.63 |
-| Vanilla | 0.30 | 0.37 | 0.49 | 0.36 | 0.42 | 0.26 | 0.35 | |
-| Janus-Pro-7B | ImAgent | 0.44 | 0.46 | 0.55 | 0.45 | 0.50 | 0.27 | 0.44 |
-| Idiom | Textual | Entity | Scientific | Overall | |||||||
-| Types | Model | Acc. | Qual. | Acc. | Qual. | Acc. | Qual. | Acc. | Qual. | Acc. | Qual. |
+| GPT-4o | 0.81 | 0.71 | 0.89 | 0.83 | 0.79 | 0.74 | 0.80 | |
+| Bagel | Vanilla | 0.44 | 0.55 | 0.68 | 0.44 | 0.60 | 0.39 | 0.52 |
+| Self-Rewrite | 0.76 | 0.69 | 0.75 | 0.65 | 0.75 | 0.58 | 0.70 | |
+| ImAgent | 0.63 | 0.63 | 0.72 | 0.59 | 0.69 | 0.53 | 0.63 | |
+| Janus-Pro-7B | Vanilla | 0.30 | 0.37 | 0.49 | 0.36 | 0.42 | 0.26 | 0.35 |
+| Self-Rewrite | 0.21 | 0.24 | 0.26 | 0.18 | 0.20 | 0.16 | 0.21 | |
+| ImAgent | 0.44 | 0.46 | 0.55 | 0.45 | 0.50 | 0.27 | 0.44 |
+| Types | Model | Idiom | Textual | Entity | Scientific | Overall | |||||
+| Acc. | Qual. | Acc. | Qual. | Acc. | Qual. | Acc. | Qual. | Acc. | Qual. | ||
 | FLUX.1-dev | 39.1 | 83.4 | 56.9 | 76.5 | 45.1 | 90.6 | 46.7 | 80.9 | 47.0 | 82.8 | |
@@
 | Gen | gpt-image-1 | 75.7 | 94.5 | 86.9 | 97.6 | 77.5 | 96.6 | 74.7 | 94.3 | 78.7 | 95.8 |
-| Emu3 | 33.1 | 82.9 | 33.7 | 68.7 | 33.8 | 85.2 | 40.1 | 77.0 | 35.2 | 78.5 | |
+| Unified | Emu3 | 33.1 | 82.9 | 33.7 | 68.7 | 33.8 | 85.2 | 40.1 | 77.0 | 35.2 | 78.5 |
 | Janus-Pro-7B | 25.5 | 78.0 | 37.2 | 70.9 | 38.5 | 87.6 | 44.9 | 77.8 | 36.5 | 78.6 | |
@@
 | GoT | 29.7 | 76.4 | 30.6 | 70.7 | 31.0 | 86.2 | 36.8 | 76.3 | 32.0 | 77.4 | |
-| Unified | Gemini-2.0 | 52.4 | 87.8 | 73.0 | 83.3 | 67.0 | 94.3 | 66.7 | 89.3 | 64.8 | 88.7 |
-| Vanilla | 30.2 | 85.7 | 36.6 | 68.4 | 45.0 | 94.7 | 54.4 | 87.5 | 41.6 | 84.1 | |
-| Bagel | ImAgent | 37.7 | 90.0 | 54.2 | 79.1 | 52.6 | 96.6 | 61.2 | 90.3 | 51.4 | 89.0 |
-| Vanilla | 25.5 | 78.0 | 37.2 | 70.9 | 38.5 | 87.6 | 44.9 | 77.8 | 36.5 | 78.6 | |
-| Janus-Pro-7B | ImAgent | 27.9 | 86.0 | 35.3 | 68.8 | 40.7 | 89.7 | 51.2 | 84.2 | 38.8 | 82.2 |
-4 Experiments
-4.1 Experimental Setting
-Models.
-We build our ImAgent on two of the most recent and powerful unified models, Bagel [deng2025emerging] and Janus-Pro-7B [chen2025janus].
-Both models possess versatile multimodal capabilities, including image understanding, image generation, which collectively form the foundation of our framework.
-For image generation, we leverage both Bagel and Janus-Pro-7B.
-However, since Janus-Pro-7B does not support image editing, we employ Bagel exclusively for the image editing tasks.
-Benchmarks.
-We evaluate ImAgent on multiple benchmarks, including 3 benchmarks on image generation and 4 benchmarks for image editing tasks.
-For image generation, we use R2I-Bench [chen2025r2i], which is designed to rigorously assess reasoning-driven T2I generation; WISE [niu2025wise], which covers six categories of image generation scenarios; and T2I-ReasonBench [sun2025t2i], which assesses the models’ reasoning ability in generative tasks.
-For image editing, we adopt GEdit-Bench [liu2504step1x], which contains both Chinese and English instruction-based editing tasks; RISEBench [zhao2025envisioning], which focuses on reasoning-informed visual editing across diverse reasoning types; KRIS-Bench [wu2025kris], which evaluates reasoning capabilities over factual, conceptual, and procedural knowledge; and ImgEdit-Bench [ye2025imgedit], which is used to evaluate image editing performance in terms of instruction adherence, editing quality, and detail preservation.
-Baselines.
-For image generation, we employ SDv1.5 [rombach2022high], SD3-Medium [rombach2022high], SDXL [podell2023sdxl], SD3.5-Medium [esser2024scaling], SD3.5-Large [esser2024scaling], PixArt-Alpha [chen2024pixart], FLUX.1-Dev [blackforest2024flux], Sana-1.5 [xie2025sana], Lumina-T2I [qin2025lumina], LLM4GENSD1.5 [liu2025llm4gen], ELLASD1.5 [hu2024ella], LlamaGen [sun2024autoregressive], DALL-E-3 [ma2025learning], gpt-image-1 [hurst2024gpt], Omnigen [xiao2025omnigen], FLUX.1-schnell [flux2024], andPlayground-v2.5 [li2024playground], as generation-based models.
-We also include Janus [wu2025janus], VILA-U [wu2024vila], Show-o [xie2024show], Janus-Pro-7B [chen2025janus], Emu3 [wang2024emu3], Lumina-Image 2.0 [qin2025lumina], show-o [xie2024show], GoT [fang2025got] and MetaQuery-XL [pan2025transfer] as unified multimodal models.
-| Types | Model | Temporal | Causal | Spatial | Logical | Overall |
-| Gemini-2.0-Flash-pre | 10.6 | 13.3 | 11.0 | 2.3 | 9.4 | |
-| Gemini-2.0 | 8.2 | 15.5 | 23.0 | 4.7 | 13.3 | |
-| Private | GPT-4o | 34.1 | 32.2 | 37.0 | 10.6 | 28.9 |
-| EMU2 | 1.2 | 1.1 | 0.0 | 0.0 | 0.5 | |
-| OmniGen | 1.2 | 1.0 | 0.0 | 1.2 | 0.8 | |
-| Step1X-Edit | 0.0 | 2.2 | 2.0 | 3.5 | 1.9 | |
-| Qwen-Image-Edit | 4.7 | 10.0 | 17.0 | 2.4 | 8.9 | |
-| FLUX.1-Kontext-Dev | 2.3 | 5.5 | 13.0 | 1.2 | 5.8 | |
-| Ovis-U1 | 1.2 | 3.3 | 4.0 | 2.4 | 2.8 | |
-| Open | Seedream-4.0 | 12.9 | 12.2 | 11.0 | 7.1 | 10.8 |
-| Vanilla | 2.4 | 5.6 | 14.0 | 1.2 | 6.1 | |
-| Bagel | ImAgent | 17.6 | 15.6 | 16.0 | 2.4 | 13.1 |
-For image editing, we evaluate both private and open baselines.
-The private baselines include Gemini-2.0 [google2025_gemini2_image], Doubao [shi2411seededit], GPT-4o [openai2025gpt4o_image], and Gemini-2.0-Flash-pre [team2023gemini].
-The open baselines include Step1X-Edit [liu2025step1x], Instruct-Pix2Pix [brooks2023instructpix2pix], MagicBrush [zhang2023magicbrush], AnyEdit [yu2025anyedit], OmniGen [xiao2025omnigen], EMU2 [sun2024generative], Qwen-Image-Edit [wu2025qwenimagetechnicalreport], FLUX.1-Kontext-Dev [labs2025flux1kontextflowmatching], Ovis-U1 [wang2025ovisu1], and Seedream-4.0 [seedream2509seedream].
-4.2 Quantitative Results
+| Gemini-2.0 | 52.4 | 87.8 | 73.0 | 83.3 | 67.0 | 94.3 | 66.7 | 89.3 | 64.8 | 88.7 | |
+| Bagel | Vanilla | 30.2 | 85.7 | 36.6 | 68.4 | 45.0 | 94.7 | 54.4 | 87.5 | 41.6 | 84.1 |
+| Self-Rewrite | 44.6 | 84.3 | 44.0 | 73.7 | 52.4 | 91.6 | 57.7 | 88.3 | 49.7 | 84.5 | |
+| ImAgent | 37.7 | 90.0 | 54.2 | 79.1 | 52.6 | 96.6 | 61.2 | 90.3 | 51.4 | 89.0 | |
+| Janus-Pro-7B | Vanilla | 25.5 | 78.0 | 37.2 | 70.9 | 38.5 | 87.6 | 44.9 | 77.8 | 36.5 | 78.6 |
+| Self-Rewrite | 11.9 | 74.6 | 9.2 | 49.3 | 12.0 | 74.3 | 14.0 | 59.3 | 11.1 | 63.3 | |
+| ImAgent | 27.9 | 86.0 | 35.3 | 68.8 | 40.7 | 89.7 | 51.2 | 84.2 | 38.8 | 82.2 | |
+| Types | Models | Add | Adjust | Extract | Replace | Remove | Background | Style | Hybrid | Action | Overall |
+| Private | GPT-4o-Image | 4.61 | 4.33 | 2.90 | 4.35 | 3.66 | 4.57 | 4.93 | 3.96 | 4.89 | 4.20 |
+| Open | MagicBrush | 2.84 | 1.58 | 1.51 | 1.97 | 1.58 | 1.75 | 2.38 | 1.62 | 1.22 | 1.83 |
+| Instruct-P2P | 2.45 | 1.83 | 1.44 | 2.01 | 1.50 | 1.44 | 3.55 | 1.20 | 1.46 | 1.88 | |
+| AnyEdit | 3.18 | 2.95 | 1.88 | 2.47 | 2.23 | 2.24 | 2.85 | 1.56 | 2.65 | 2.45 | |
+| UltraEdit | 3.44 | 2.81 | 2.13 | 2.96 | 1.45 | 2.83 | 3.76 | 1.91 | 2.98 | 2.70 | |
+| Step1X-Edit | 3.88 | 3.14 | 1.76 | 3.40 | 2.41 | 3.16 | 4.63 | 2.64 | 2.52 | 3.06 | |
+| UniWorld-V1 | 3.82 | 3.64 | 2.27 | 3.47 | 3.24 | 2.99 | 4.21 | 2.96 | 2.74 | 3.26 | |
+| Bagel | Vanilla | 3.89 | 3.68 | 1.33 | 3.78 | 3.07 | 2.07 | 4.30 | 2.47 | 4.33 | 2.89 |
+| Self-Rewrite | 3.22 | 3.25 | 2.31 | 3.55 | 2.13 | 3.41 | 3.97 | 2.61 | 4.56 | 3.00 | |
+| ImAgent | 3.20 | 3.89 | 1.77 | 3.13 | 2.71 | 3.50 | 4.21 | 2.89 | 4.43 | 3.15 |
+| Types | Models | EN Setting | CN Setting | ||||
+| G_SC | G_PQ | G_O | G_SC | G_PQ | G_O | ||
+| Private | Gemini-2.0 | 6.73 | 6.61 | 6.32 | 5.43 | 6.77 | 5.36 |
+| Doubao | 6.92 | 7.19 | 6.75 | 6.98 | 7.27 | 6.77 | |
+| GPT-4o | 7.85 | 7.62 | 7.53 | 7.67 | 7.56 | 7.30 | |
+| Open | Instruct-Pix2Pix | 3.58 | 5.50 | 3.68 | - | - | - |
+| MagicBrush | 4.68 | 5.66 | 4.52 | - | - | - | |
+| AnyEdit | 3.18 | 5.82 | 3.21 | - | - | - | |
+| OmniGen | 5.96 | 5.89 | 5.06 | - | - | - | |
+| Step1X-Edit | 7.09 | 6.76 | 6.70 | 7.20 | 6.87 | 6.86 | |
+| Bagel | Vanilla | 7.36 | 6.83 | 6.52 | 7.34 | 6.85 | 6.50 |
+| Self-Rewrite | 7.66 | 6.57 | 6.61 | 7.50 | 6.60 | 6.58 | |
+| ImAgent | 7.85 | 6.61 | 6.88 | 7.92 | 6.53 | 6.84 | |
+4.2 Main Results
 Image Generation.
@@
 Compared to their respective backbone models, ImAgent built upon Bagel and Janus-Pro-7B consistently surpasses the vanilla counterparts, with performance improvements of 21.2% and 25.7%, respectively, demonstrating the effectiveness of our approach in enhancing image generation quality through coordinated reasoning and refinement.
-Notably, the vanilla Janus-Pro-7B underperforms compared to SDXL and EMU3 and performs on par with Show-o.
+Furthermore, the vanilla Janus-Pro-7B underperforms compared to SDXL and EMU3 and performs on par with Show-o.
 In contrast, ImAgent constructed on Janus-Pro-7B surpasses all of them, highlighting its strong test-time scaling capability and generalization potential.
 Furthermore, on the reasoning benchmarks, ImAgent consistently delivers strong results.
-As shown in Tables 1 and 3, ImAgent outperforms the vanilla models across both backbone architectures.
+As shown in Tables 1 and Table 4, ImAgent outperforms the vanilla models across both backbone architectures.
 For instance, ImAgent achieves improvements of 14.8% and 7.5% across both backbones on R2I-Bench, respectively.
 Moreover, ImAgent built upon Bagel surpasses the commercial model Gemini-2.0 in overall quality on the T2I-ReasonBench, further demonstrating its effectiveness in reasoning-based image generation.
+Notably, ImAgent constructed on Bagel underperforms the Self-Rewrite baseline on WISE.
+To further investigate this phenomenon, we perform an additional control experiment by rewriting the prompts using GPT-4o with the same system prompt as used in the Bagel Self-Rewrite implementation.
+Surprisingly, this GPT-4o-based rewrite achieves a score of 0.64, which is significantly lower than the Self-Rewrite performance.
+This suggests that the strong performance of Self-Rewrite on this dataset may benefit from dataset-specific familiarity.
+Additionally, Self-Rewrite underperforms Vanilla on Janus-Pro-7B, whereas ImAgent constructed on Janus-Pro-7B consistently outperforms Vanilla, using the same prompt as in the Prompt Enhancement with CoT action.
+This observation suggests that Janus-Pro-7B is less effective in text-only prompt rewriting, while ImAgent mitigates this limitation by leveraging test-time agentic control and scalable decision-making.
 Image Editing.
@@
 This indicates that open-source models enhanced with ImAgent can achieve competitive performance with commercial models.
-Furthermore, as shown in Figure 4, ImAgent also achieves excellent performance on KRISBench, attaining an overall score of 67.13, surpassing vanilla Bagel which scores 63.16.
-ImAgent performs well on ImgEdit-Bench (shown in Figure 3) as well, with vanilla achieving 2.89 and underperforming compared to Step1X-Edit, while ImAgent reaches 3.15, outperforming Step1X-Edit and achieving the best performance among open-source models.
+Furthermore, ImAgent performs well on ImgEdit-Bench (shown in Table 5) as well, with vanilla achieving 2.89 and underperforming compared to Step1X-Edit, while ImAgent reaches 3.15, outperforming Step1X-Edit and achieving the best performance among open-source models.
 This further demonstrates the effectiveness of ImAgent in test-time scaling.
-Moreover, ImAgent outperforms vanilla Bagel across both English and Chinese settings in GEdit-Bench, achieving improvements of 5.5% and 5.2%, respectively.
+Moreover, as shown in Table 6, ImAgent outperforms vanilla Bagel across both English and Chinese settings in GEdit-Bench, achieving improvements of 5.5% and 5.2%, respectively.
 Detailed results are provided in the Appendix.
+4.3 Robustness Analysis
+Beyond the strong quantitative results across various benchmarks, robustness is critical for the reliable deployment and generalization of training-free architectures.
+We therefore conduct a robustness analysis of ImAgent using three metrics: Parse Success Rate (PSR), Action Uniqueness Rate (AUR), and Action Validity Rate (AVR).
+PSR measures whether the model successfully parses an action, AUR evaluates whether the model outputs a single valid action rather than hallucinating multiple actions, and AVR measures the probability that the predicted action belongs to the predefined action set.
+The detailed definitions of these metrics are available in Appendix.
+[…]
+| Dataset | PSR (%) | AUR (%) | AVR (%) |
+| WISEBagel | 100% | 100% | 100% |
+| WISEJanus | 99% | 100% | 100% |
+| RISEBenchBagel | 100% | 100% | 100% |
+4.4 Agentic Behavior Analysis
+We compare the agentic behaviors of ImAgent constructed on different unified backbone models by analyzing their decision trajectories. We derive two key observations.
+Observation 1: The average decision length differs across backbone models.
+Specifically, ImAgent built on Bagel consumes more computational budget than Janus-Pro-7B.
+Under the image generation task, ImAgent on Bagel executes an average of 4.73 iterations, whereas ImAgent on Janus-Pro-7B requires only 3.98 iterations on average.
+This result indicates that, even with identical workflows and prompts, different unified backbones exhibit distinct agentic behaviors and computational resource utilization patterns.
+We hypothesize that Bagel tends to perform more conservative or iterative refinement, while Janus-Pro-7B exhibits more decisive behavior, leading to fewer interaction steps.
+Observation 2: Both backbone models exhibit similar action selection trends.
+In particular, they tend to invoke the Best-of-N Sampling action at later stages of the interaction, unless computational constraints prevent its execution, as the policy controller recognizes that this action incurs higher computational cost.
+Moreover, ImAgent invokes the Image Detail Refinement action less frequently than other actions.
+This behavior is intuitive, as this action requires the model to first generate a localized editing prompt conditioned on both the generated image and the user prompt, and then perform image editing, which introduces additional complexity and uncertainty.
+This suggests that the ImAgent implicitly prefers lower-cost and lower-risk actions, resembling a rational cost–benefit trade-off strategy.
+4.5 Test-Time Scalability Analysis.
+To demonstrate the test-time scalability of ImAgent, we compare it with a strong training-free iterative refinement baseline under the same inference budget.
+Specifically, we implement an iterative prompt self-refinement strategy, where the unified model repeatedly rewrites the prompt based on the generated image and regenerates the image for steps, matching the number of model calls used by ImAgent.
+As shown in Table 8, ImAgent consistently outperforms iterative refinement under the same compute budget.
+These results indicate that adaptive policy-driven test-time scaling is more effective than naive iterative refinement.
+| Method | R2I-Bench | T2I (Acc.) | T2I (Qual.) |
+| Vanilla | 0.54 | 41.6 | 84.1 |
+| Self-Rewrite | 0.58 | 49.7 | 84.5 |
+| Iterative-Refinement | 0.60 | 50.3 | 86.6 |
+| ImAgent | 0.62 | 51.4 | 89.0 |
+| Benchmark | action1 | action2 | action3 | action4 | action5 | Random | ImAgent |
+| R2I-Bench | 0.54 | 0.58 | 0.57 | 0.56 | 0.56 | 0.59 | 0.62 |
+| T2I (Acc.) | 41.6 | 49.7 | 45.3 | 43.5 | 46.2 | 44.2 | 51.4 |
+| T2I (Qual.) | 84.1 | 84.5 | 84.2 | 86.4 | 85.2 | 87.4 | 89.0 |
 4.6 Qualitative Analysis
-Image Refinement Flow.
-As shown in Figure 1, we present several qualitative examples, including two for image generation and two for image editing, each executed for five steps.
-From an overall perspective, we can observe that the image quality improves progressively with each step, demonstrating that ImAgent refines the image gradually.
-Specifically, consider a case where the user aims to generate an image based on the prompt “moldy bread.” In the first step, ImAgent performs the Naive Generation action and produces an image containing fresh bread, which clearly deviates from the prompt.
-The policy controller then selects the Prompt Refinement action to improve the prompt using both the generated image and the user’s original input.
-Although the updated image shows slight improvement, it remains unsatisfactory.
-To address this, the controller chooses the Image Detail Refinement action, leading to the appearance of mold on the bread.
-Subsequently, ImAgent applies the Prompt Enhancement action to further strengthen the textual guidance, resulting in an image that better aligns with the user’s intent.
-Finally, recognizing that the prompt is now sufficiently clear but the visual result could still be improved, ImAgent executes the Best-of-N Sampling action to generate multiple candidates and select the most aligned one.
-The final output successfully captures the desired “moldy bread” concept.
-Detailed Policy Controller Decision.
-We present a detailed case study of the policy controller’s output to illustrate how it makes action decisions.
-As shown in Figure 6, which visualizes the decision-making process for the case discussed earlier, the policy controller selects the promptrefinement action to enrich the textual prompt and improve semantic clarity for subsequent generation steps.
-We present more examples in the Appendix.
-4.4 Ablation Studies
-Comparison with Self-CoT.
-Since the vanilla Bagel model already incorporates self-CoT capabilities, we further compare ImAgent with self-CoT to evaluate the effectiveness of our approach.
-For the image editing task, as shown in Figure 5(a), ImAgent significantly outperforms both the vanilla and self-CoT variants on the Gedit-Bench across both English and Chinese settings.
-Furthermore, ImAgent also surpasses self-CoT in the image generation task.
-As illustrated in Figures 5(b) and (c), under the Janus-Pro-7B backbone, the performance of self-CoT notably drops compared to the vanilla model, whereas ImAgent consistently achieves superior results.
-Similarly, under the Bagel backbone, ImAgent maintains clear advantages over both the vanilla and self-CoT models.
-These findings suggest that the improvements brought by ImAgent stem not only from self-CoT–like reasoning refinement, but also from the different actions coordination enabled by the strong policy controller.
-Effectiveness of the Whole ImAgent.
+As shown in Figure LABEL:fig:example, we provide several qualitative examples, including two for image generation and two for image editing, each executed over five iterative steps.
+Overall, we observe that the image quality improves progressively at each step, demonstrating that ImAgent gradually refines the output.
+Specifically, consider a case where the user intends to generate an image based on the prompt “moldy bread.” Here, ImAgent dynamically selects a sequence of five actions to generate the final image.
+The complete decision history for this case is provided in Appendix 8.
+To illustrate the policy controller’s reasoning, Figure 4 visualizes how the controller selects actions at each step.
+For instance, it chooses the promptrefinement action to enrich the textual prompt, thereby improving semantic clarity for subsequent generation steps.
+Additional examples of policy decisions are presented in Appendix 8.
+4.7 Effectiveness of the Whole ImAgent.
 To evaluate the effectiveness of ImAgent in coordinately selecting and executing appropriate actions, we compare its performance with that of single-turn actions and a random action selection policy, where the for random selection is set the same as in ImAgent.
-As shown in Table 5, experiments are conducted on R2I-Bench [chen2025r2i].
-All single-turn actions perform well compared to the naive generation action, which serves as the lower bound (0.54).
+As shown in Table 9, experiments are conducted on R2I-Bench and T2I-ReasonBench.
+Taking R2I-Bench as an example, all single-turn actions perform well compared to the naive generation action, which serves as the lower bound (0.54).
 Moreover, the random action selection policy achieves a score of 0.59, higher than all single-turn actions, indicating that naively constructing an agent is beneficial.
 Finally, ImAgent outperforms the random policy with a score of 0.62, demonstrating that it effectively and efficiently boosts test-time scaling in image generation.
-| Method | Random | ImAgent | |||||
-| Score | 0.54 | 0.58 | 0.57 | 0.56 | 0.56 | 0.59 | 0.62 |
 5 Conclusion
@@
 Supplementary Material
-Appendix A Algorithm for Image Editing
+6 Algorithm for Image Editing
 We present the detailed algorithm (shown in Algorithm 2) for image editing.
-Appendix B Detailed Action Description
+7 Detailed Action Description
 Naive Generation/Editing.
@@
 By exposing the implicit reasoning steps, the model produces outputs that are more semantically aligned, compositionally consistent, and visually faithful to user intent.
-| Method | Factual | Conceptual | Procedural | Overall | |||
-| AP | SP | SS | NS | LP | ID | ||
-| Vanilla | 73.27 | 67.17 | 61.10 | 62.36 | 54.38 | 64.77 | 63.16 |
-| ImAgent | 74.39 | 71.15 | 69.80 | 66.03 | 54.83 | 68.78 | 67.13 |
-| Types | Models | Add | Adjust | Extract | Replace | Remove | Background | Style | Hybrid | Action | Overall |
-| Private | GPT-4o-Image | 4.61 | 4.33 | 2.90 | 4.35 | 3.66 | 4.57 | 4.93 | 3.96 | 4.89 | 4.20 |
-| Open | MagicBrush | 2.84 | 1.58 | 1.51 | 1.97 | 1.58 | 1.75 | 2.38 | 1.62 | 1.22 | 1.83 |
-| Instruct-P2P | 2.45 | 1.83 | 1.44 | 2.01 | 1.50 | 1.44 | 3.55 | 1.20 | 1.46 | 1.88 | |
-| AnyEdit | 3.18 | 2.95 | 1.88 | 2.47 | 2.23 | 2.24 | 2.85 | 1.56 | 2.65 | 2.45 | |
-| UltraEdit | 3.44 | 2.81 | 2.13 | 2.96 | 1.45 | 2.83 | 3.76 | 1.91 | 2.98 | 2.70 | |
-| Step1X-Edit | 3.88 | 3.14 | 1.76 | 3.40 | 2.41 | 3.16 | 4.63 | 2.64 | 2.52 | 3.06 | |
-| UniWorld-V1 | 3.82 | 3.64 | 2.27 | 3.47 | 3.24 | 2.99 | 4.21 | 2.96 | 2.74 | 3.26 | |
-| Bagel | Vanilla | 3.89 | 3.68 | 1.33 | 3.78 | 3.07 | 2.07 | 4.30 | 2.47 | 4.33 | 2.89 |
-| ImAgent | 3.20 | 3.89 | 1.77 | 3.13 | 2.71 | 3.50 | 4.21 | 2.89 | 4.43 | 3.15 |
 Prompt Revision Based on the Generated/Edited Image.
@@
 This sampling strategy stabilizes generation quality and significantly mitigates randomness caused by diffusion or autoregressive sampling.
-Appendix C Detailed Experimental Results
-Appendix D Detailed Policy Controller Decision
+8 Detailed Policy Controller Decision
 We present a case study that illustrates the step-by-step decisions made by the policy controller.
@@
 For the fifth step, ImAgent decides to invoke the Best-of-N Sampling:
-| Types | Models | EN Setting | CN Setting | ||||
-| G_SC | G_PQ | G_O | G_SC | G_PQ | G_O | ||
-| Private | Gemini-2.0 | 6.73 | 6.61 | 6.32 | 5.43 | 6.77 | 5.36 |
-| Doubao | 6.92 | 7.19 | 6.75 | 6.98 | 7.27 | 6.77 | |
-| GPT-4o | 7.85 | 7.62 | 7.53 | 7.67 | 7.56 | 7.30 | |
-| Open | Instruct-Pix2Pix | 3.58 | 5.50 | 3.68 | - | - | - |
-| MagicBrush | 4.68 | 5.66 | 4.52 | - | - | - | |
-| AnyEdit | 3.18 | 5.82 | 3.21 | - | - | - | |
-| OmniGen | 5.96 | 5.89 | 5.06 | - | - | - | |
-| Step1X-Edit | 7.09 | 6.76 | 6.70 | 7.20 | 6.87 | 6.86 | |
-| Bagel | Vanilla | 7.36 | 6.83 | 6.52 | 7.34 | 6.85 | 6.50 |
-| ImAgent | 7.85 | 6.61 | 6.88 | 7.92 | 6.53 | 6.84 | |
+[…]
+9 Detailed Prompts for Each Action.
+We present the detailed prompts used for each action in ImAgent.
+Specifically, Figure 5 presents the prompt for Prompt Enhancement with CoT, and Figure 6 provides the prompt for Prompt Revision Based on the Generated/Edited Image.
+Figure 7 shows the prompt for Image Detail Refinement, while Figure 8 presents the evaluation prompt used in action Best-of-N Sampling.
+The actions Naive Generation/Editing and STOP directly use the current prompt and therefore do not require additional prompting templates.